OpenStudy (anonymous):

a 20 ft ladder is placed 12 feet away from the post. how high is the post

OpenStudy (anonymous):

pls fast this answer i need it tomorrow :(

OpenStudy (anonymous):



OpenStudy (anonymous):

When you can sketch, why not?

OpenStudy (sheraz12345):

Apply Pythagoras therom

OpenStudy (anonymous):

With the post perpendicular to the floor you can see a 90 degree angle is formed.

OpenStudy (anonymous):

Pythagoras' theorem states that a^2 = b^2 + c^2 where a is the hypotenuse (in this case the ladder is the hypotenuse which is easiest to surmise when you realise that the ladder will be opposite the ninety degree angle created by the flat earth and the post perpendicular to it), and b and c are the other sides of the right-angled triangle. We have the length of a and either b or c (doesn't matter which) so 20^2 = 12^2 + c^2 I.e. c^2 = 400 - 144 c^2 = 256 c = 16 This means the length of the post is 16 feet tall

OpenStudy (anonymous):

Oops, sorry! The length of the post is 16 feet; the length isn't 16 feet tall :P

OpenStudy (anonymous):

Did I answer your question?